OpenStudy (slinley):

Which shows the expression below simplified? 0.0000042 ÷ (7 multiplyed by 10^-4) 6 × 10^-1 6 × 10^-4 6 × 10^-3 6 × 10^-2 @Catlover5925

OpenStudy (catlover5925):

ok 10^-4 is 0.0001so multiply that by 7 and it is 0.0007 Then take 0.0000042 and divide it by 0.0007 Your final answer should be 0.006 or over left (negative) 3 places ANSWER IS C --> 6 * 10^-3

OpenStudy (slinley):

Thank you! :)

OpenStudy (catlover5925):

Np anytime!!

OpenStudy (slinley):

can you help me on another question? :D

OpenStudy (slinley):

@Catlover5925

OpenStudy (catlover5925):

yes what is it?

OpenStudy (slinley):

The contents of a cell in a spreadsheet are shown below. 1.0311889E-6 What number do the contents of the cell represent? 0.0000001031 1,031,188.9 103,118.8900000000 0.0000010312

OpenStudy (catlover5925):

Im sooo sorry I don't know.....

OpenStudy (slinley):

Ok! Its fine! :D

OpenStudy (slinley):

, Which shows the expression below simplified? (6 multiplyed by 10^4) × 0.05 3 multiplyed by 10^3 3 multiplyed by 10^2 1.1 multiplyed by 10^-8 3 multiplyed by 10^-8

OpenStudy (slinley):

those question marks are supposed to be a multiplication sign!

zepdrix (zepdrix):

0_o

zepdrix (zepdrix):

\[\Large\rm (6\times 10^4)\times 0.05\]Multiplication is `commutative`, meaning, you can multiply things in any order that you want. So multiply the 6 and 0.05, what do you get? :U

OpenStudy (slinley):

0.3 ????

zepdrix (zepdrix):

\[\Large\rm 0.3\times 10^4\]Ok good, almost done.

zepdrix (zepdrix):

In scientific notation, we want that leading term to be greater than 1, but less than 10.

zepdrix (zepdrix):

But right now it's smaller than 1 :( So we'll need to borrow a 10 from the 10^4 guy.

zepdrix (zepdrix):

\[\Large\rm 0.3\times 10\times10^3\]

zepdrix (zepdrix):

Multiplying by 10 just moves the decimal place over one, yes?

OpenStudy (slinley):

Right,, now i've have learned something newww

OpenStudy (slinley):

thx

zepdrix (zepdrix):

Figure out which one it is? c:

OpenStudy (slinley):

yes, think so

OpenStudy (slinley):

6*10^-3

zepdrix (zepdrix):

Oh the previous problem: `The contents of a cell in a spreadsheet are shown below.` `1.0311889E-6` E-6 is just a short way for the calculator/computer to write \(\Large\rm \times 10^{-6}\)

zepdrix (zepdrix):

Sorry I gotta get ready for school :c

OpenStudy (slinley):

ok lol ur now going to school? Its 11:29 am!